OpenStudy (tanabugg):

-3m + 3 (m + 6) 6 -6m + 6 6m + 18 18

OpenStudy (anonymous):

18

OpenStudy (tanabugg):

You are a life saver thanks!

OpenStudy (anonymous):

sure! do you need an explanation?

OpenStudy (tanabugg):

Um yes please.

OpenStudy (anonymous):

-3m+3(m+6) 3m+6 -3m+3m+18 -3m+3m=0+18=18

OpenStudy (anonymous):

The key idea is to distribute the '3' outside of the parentheses

OpenStudy (tanabugg):

Okay I got it! Thanks so much

OpenStudy (anonymous):

Ya welcome